6. [ 10 pts.] Let X be a random variable with and σ,= 4. Find the following quantities: a. E[2X - 4] b. E[X2] c. E[4X2]
Answer #1

Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

E[X]=1 and Var(X)=4

So,

a) E[2X-4]=2*E[X]-4=2-4=-2

b) var(X)=E[X^2]-E[X]^2

So,

E[X^2]=4-1=3

c) E[4*X^2]=4*E[X^2]=4*3=12

d) E[(2X-4)^2]=E[4X^2+16-16X]=4*E[X^2]+16-16*E[X]=4*3+16-16*1=12

Kindly revert for any queries

Thanks.